# Generative and Protective AI for Content Creation

### v1

@@ Appendix C Details on Dataset Inference
Diffusion models. For DMs, the CDI methodology [dubinski2025cdi] employs a broad feature set, which we describe in Appendix B. Rather than aggregating these features directly, CDI fits a logistic regression scorer on disjoint control splits of $P$ and $U$, yielding a calibrated mapping from feature vectors to scalar scores. The test is then applied to scores on held-out subsets. To reduce variance, CDI repeats this process across multiple random partitions and averages the resulting $p$-values.
Image autoregressive models. For IARs, the approach of [kowalczukprivacy] follows the same overall structure but makes use of a different feature suite, tailored to token-level modeling. These features, described in Appendix A, capture variations in token probabilities and loss trajectories that arise in autoregressive generation. Each feature is normalized, and the per-sample scalar score is obtained by summing across all features. This lighter-weight procedure that nevertheless suffices in practice for autoregressive token-based models. The resulting scores for $P$ and $U$ are then compared using the same statistical test as above.

### v2

@@ Appendix C Details on Dataset Inference
Diffusion models. For DMs, the CDI methodology [7] employs a broad feature set, which we describe in Appendix B. Rather than aggregating these features directly, CDI fits a logistic regression scorer on disjoint control splits of $P$ and $U$, yielding a calibrated mapping from feature vectors to scalar scores. The test is then applied to scores on held-out subsets. To reduce variance, CDI repeats this process across multiple random partitions and averages the resulting $p$-values.
Image autoregressive models. For IARs, the approach of [17] follows the same overall structure but makes use of a different feature suite, tailored to token-level modeling. These features, described in Appendix A, capture variations in token probabilities and loss trajectories that arise in autoregressive generation. Each feature is normalized, and the per-sample scalar score is obtained by summing across all features. This lighter-weight procedure that nevertheless suffices in practice for autoregressive token-based models. The resulting scores for $P$ and $U$ are then compared using the same statistical test as above.